OpenStudy (anonymous):

Write the equation x^2 + 7y^2 - 12x - 13 in standard form. State which type of conic section it is. Please Help Medals will be given.

OpenStudy (anonymous):

(x^2-12x+____)+7(y^2)=13+____ (-12/2)^2=(-6)^2=36 (x^2-12x+36)+7(y^2)=13+36 (x-6)^2+(y)^2=49 I believe i might be missing a step but the center of this circle is at (6,0) with a radius of 7. Let me know if this helps